OpenStudy (anonymous):

let f(x)=(absolute val of x)/(x). show that f(x)=1,x>0 f(x)=-1,x<0 find the domain and range of f(x)

OpenStudy (anonymous):

domain all numbers except 0

OpenStudy (anonymous):

range is 1 and -1

OpenStudy (anonymous):

i kind of have to explain how i know that though

OpenStudy (anonymous):

\[f(x) = \frac{|x|}{x} = \left\{\begin{array}{rcc} 1 & \text{if} & x >0 \\ -1& \text{if} & x < 0 \end{array} \right.\]

OpenStudy (anonymous):

Well, if x > 0 then |x| = x Therefore |x|/x = x/x = 1 If x < 0 then |x| = -x Therefore |x|/x = -x/x = -1

OpenStudy (anonymous):

ok well we start with \[ |x| = \left\{\begin{array}{rcc} x & \text{if} & x \geq 0 \\ - x& \text{if} & x < 0 \end{array} \right.\]

OpenStudy (anonymous):

that is the definition. then divide by x to get what i wrote above exactly.

OpenStudy (anonymous):

what polpak said. hello polpak

OpenStudy (anonymous):

Nice latex btw =)

OpenStudy (anonymous):

merci. i stole it from zarkon

OpenStudy (anonymous):

And I shall steal it from you ;p Wouldn't have thought to lay that out using an array